OpenStudy (anonymous):

which fraction is equivalent to5 3/4? A. 19/4 B. 23/4 C. 8/4 D. 15/4

jimthompson5910 (jim_thompson5910):

I'm assuming you mean \[\Large 5 \frac{3}{4}\]

jimthompson5910 (jim_thompson5910):

it might help to see it as \[\Large 5 \frac{3}{4} = 5 + \frac{3}{4}\]

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

now multiply top and bottom of 5 by 4 to get a common denominator \[\Large 5 \frac{3}{4} = 5 + \frac{3}{4} = 5{\color{red}{*\frac{4}{4}}} + \frac{3}{4}\]

OpenStudy (anonymous):

12

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

\[\Large 5 \frac{3}{4} = 5 + \frac{3}{4} = 5{\color{red}{*\frac{4}{4}}} + \frac{3}{4} = \frac{5*4}{4} + \frac{3}{4} = ??\]

OpenStudy (anonymous):

23

jimthompson5910 (jim_thompson5910):

23/4, yes

OpenStudy (anonymous):

b